OpenStudy (nikki_skellington):

Part A: Is the probability of hitting the black circle inside the target closer to 0 or 1? Explain your answer and show your work. Part B: Is the probability of hitting the white portion of the target closer to 0 or 1? Explain your answer and show your work.

OpenStudy (nikki_skellington):

Sophia throws a dart at this square-shaped target:

OpenStudy (xguardians):

Fine the area of the circle and the area of the entire rectangle.

OpenStudy (nikki_skellington):

but i am not sure on how to find the area of a circle...

OpenStudy (anonymous):

diameter=3 radius=3/2=1.5 area of circle= pi r^2=pi*(1.5)^2=?

OpenStudy (nikki_skellington):

so the area of the circle is 1.5?